The first term of the sequence is -2 and the fifth term is -1/8. List 4 things you know. (Recursive formula, explicit formula, missing terms in sequence, common difference or ratio, whether it's geometric or arithmetic)
Answer by richwmiller(17219)

-1/8 =-2 + (4)d
d=15/32
an=a(n-1)+15/32
an = -2 + 15/32(n - 1)
an=-2+15/32n-15/32

-1/8= -2 * r^(4)
r=1/2 or -1/2
Tn = t * r^(n - 1)
It could be geometric or it could be arithmetic